For the following utility functions, a. Find the marginal rate of substitution. b. Derive the equation...

  1. For the following utility functions,

a. Find the marginal rate of substitution.


b. Derive the equation for the indifference curve where utility is equal to a value of 100.

c. Plot the indifference curve where utility is equal to a value of 100.


(1) u(x1, x2) = x1x2;

(2) u(x1, x2) = x1x2 + 10x2;

(3) u(x1, x2) = x12 + x2
Answer #1


(1) u(x1, x2) = x1x2;

a. MRS=MUx1/MUx2= x2/x1

b. 100=x1x2

c. 25 X2 20 -15 10 5 0 5 10 15 x1

(2) u(x1, x2) = x1x2 + 10x2;

a. MRS=MUx1/MUx2=x2/(x1+10)

b. 100=x1x2+10x2

c.

III x2 15 -10 -5 0 5 10 15 x1

(3) u(x1, x2) = x12 + x2

a. MRS= MUx1/MUx2=2x1/1

b. 100=x12+x2

c. 三g 100 50 5 10 0 15 x1